Given the following production plan for a firm in the short run.  

i.At what point is the firm’s Marginal Physical Product (MPP) Maximum? (5 Marks) ii.Calculate the levels of Average Physical Product (APP) and Marginal Physical Product (MPP) at  each input level. (20 Marks)

APP 

MPP




5



11



16



20



23



25



26



26



25




1
Expert's answer
2021-06-03T13:51:04-0400

Solution:

i.). The firm’s Marginal Physical Product (MPP) is maximum at the inflection point. At this point the incremental unit of the input X is at its greatest. At the inflection point, MPP equals to zero and TPP reaches its maximum, that is output is maximized at that particular point.

 

ii). Marginal Physical Product (MPP) is the change in the level of output (TPP or Y) due to a change in the variable input (X).


MPP = "\frac{\triangle TPP}{\triangle X} = \frac{TPP_{2} - TPP_{1} }{X_{2} - X_{1}}"


Average Physical Product (APP) is the quantity of output (TPP) divided by the unit of the variable input (X).


APP = "\frac{TPP}{X}"


This has been calculated in the below graph: